Your work sends 2 people on a trip,using either 2 men or 2 women. There are 8 men and 5 women. How many ways are there to do this?
Answer by ewatrrr(24785)


Hi
12 people:  8M  5F
chose two, either 2 men OR 2 women.
Ways of doing this:  8C2 + 5C2 = 28 + 10 = 38 ways

Wish You the Best in your Studies.